A heat pump picks up 1000kJ of heat from well water at 10^oc
C and discharges 3000kJ of
heat to a building to maintain it at 20^oc
C. What is the COP of the heat pump? What is the
minimum required cycle network? Please this is the right answer i need the solution. (1.5, 2000kJ)
1
Expert's answer
2020-11-02T09:25:39-0500

The minimum required cycle network


"W=Q_H-Q_C=3000-1000=2000\ kJ"

"COP=\frac{Q_H}{W}=\frac{3000}{2000}=1.5"